OpenStudy (anonymous):

Convert the equation from polar to rectangular form r^2= theta

OpenStudy (johnweldon1993):

Well \[\large r^2 = x^2 + y^2\] And \[\large \theta = \tan^{-1}\frac{y}{x}\]

OpenStudy (anonymous):

thank you

OpenStudy (johnweldon1993):

Not a problem, it would be good to remember how to derive those formulas, they can come in handy :)